40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Nov 29 14:22:03 2011, Zach, Update, RF System, EOM temp stabilization performance Screen_shot_2011-11-29_at_1.33.50_PM.pngScreen_shot_2011-11-29_at_1.34.06_PM.pngScreen_shot_2011-11-29_at_1.41.03_PM.png
    Reply  Tue Nov 29 15:25:29 2011, steve, Update, RF System, EOM temp stabilization performance 2days.png
       Reply  Tue Nov 29 18:17:27 2011, Zach, Update, RF System, EOM temp stabilization performance 
    Reply  Tue Nov 29 22:10:18 2011, kiwamu, Update, RF System, RFAM fluctuation reduced RFAM_withEOMheater_edit.png
       Reply  Tue Nov 29 23:03:34 2011, Zach, Update, RF System, RFAM fluctuation reduced 
          Reply  Wed Nov 30 03:01:55 2011, kiwamu, Update, RF System, RFAM fluctuation reduced 
Message ID: 6036     Entry time: Tue Nov 29 15:25:29 2011     In reply to: 6035     Reply to this: 6040
Author: steve 
Type: Update 
Category: RF System 
Subject: EOM temp stabilization performance 

 It is not obvious what is working.

 

Attachment 1: 2days.png  56 kB  | Hide | Hide all
2days.png
ELOG V3.1.3-